# TEMPLE: Temporal Preference Learning of Video LLMs via Difficulty Scheduling and Pre-SFT Alignment

### v2

@@ 3 Experiment
#Params VideoMME MLVU Vinoground
TP TR Overall Overall Text Video Group Average
Closed-source Models
Gemini-1.5-pro [9] - - - 75.0 - 35.8 22.6 10.2 22.9
GPT-4o [22] - - - 77.2 64.6 54.0 38.2 24.6 38.9
Claude 3.5 Sonnet [1] - - - 60.0 - 32.8 28.8 10.6 24.1
Open-source Models
Qwen2-VL [31] 72B - - 71.2 - 50.4 32.6 17.4 33.5
VideoLLaMA2 [6] 72B - - 62.4 61.2 36.2 21.8 8.4 22.1
TPO [15] 7B - - 65.6 71.1 - - - -
Tarsier 2 [34] 7B - - 64.5 67.9 - - - -
LLaVA-Video [39] 7B - - 63.3 70.8 - - - -
LLaVA-OneVision [13] 7B - - 58.2 64.7 41.6 29.4 14.6 28.5
MiniCPM-V-2.6 [5] 8B - - 60.9 - 32.6 29.2 11.2 24.3
Video-LLaVA [16] 7B - - 39.9 47.3 24.8 25.8 6.6 19.1
Qwen2-VL-2B 2B 54.5 34.5 55.2 57.7 26.4 24.6 6.2 19.1
   + SFT 2B 65.5 32.8 54.6 60.0 27.4 26.0 5.4 19.6
   + TEMPLE (Ours) 2B 63.6  (+9.1) 35.0  (+0.5) 55.1 (-0.1) 60.2  (+2.5) 28.0  (+1.6) 25.2  (+0.6) 5.2 (-1.0) 19.5  (+0.4)
Qwen2-VL-7B 7B 70.9 42.4 60.4 64.5 34.0 26.6 10.6 23.7
   + SFT 7B 70.9 44.1 60.7 64.6 34.6 26.6 9.6 23.6
   + TEMPLE (Ours) 7B 74.5  (+3.6) 45.2  (+2.8) 61.0  (+0.6) 65.8  (+1.3) 37.2  (+3.2) 29.6  (+3.0) 12.2  (+1.6) 26.3  (+2.6)
Table 1: Benchmark results of our approach on VideoMME, MLVU, and Vinoground. TP and TR denote Temporal Perception and Temporal Reasoning, respectively, two subtasks of VideoMME.
@@ 3.1 Setting
Table 2: The number of remaining videos after each step of our data pipeline.
We conduct our main experiments using the Qwen2-VL model family [31], focusing on two model scales: 2B and 7B parameters, due to its strong performance in open-source video LLMs. During data construction, the videos are sourced from a subset of the LLaVA-Video-178K dataset [39], i.e., Youtube videos within the duration of 1-3 minutes. After each step of our data pipeline, the number of remaining videos is shown in Table 2. To systematically analyze the impact of perturbation difficulty, we introduce four levels of difficulty, with the value of $r$ as 2, 4, 8, and 16. The final data contains 6,385 training samples for each difficulty level, resulting in a total of 25,540 examples.
@@
For DPO training, we adopt a four-stage progressive difficulty schedule, where the model is trained sequentially on increasingly challenging data. Beginning with the easiest perturbation setting ($r=16$), the model gradually adapts to more complex data at $r=8$, $r=4$, and finally $r=2$. At each stage, only the data corresponding to that specific difficulty is used. To ensure efficient training, we apply LoRA [10] with rank 32, $\alpha=64$, a batch size of 16, and a learning rate of 5e-6, while keeping the ViT module frozen. Each training stage lasts for approximately one hour on 8 NVIDIA H800 GPUs, accumulating to a total of four hours for the entire DPO process.
@@
For SFT, we use a dataset of 60,000 samples randomly drawn from LLaVA-Video-178K [39]. Unlike in DPO, the ViT module is unfrozen to allow for additional adaptation. The SFT phase follows a similar training setup with LoRA rank set to 32, $\alpha$ set to 32, a batch size of 16, and a higher learning rate of 1e-4. The entire fine-tuning process takes approximately three hours on the same hardware setup.
@@
Both DPO and SFT follow the original Qwen2-VL preprocessing pipeline, where videos are sampled at 2 FPS. While extracted keyframes are used for data construction, they are not used during training. To accommodate memory constraints, we limit the number of pixels per frame to 90,000, the maximum number of frames per video to 100, and the total sequence length to 8,192.
Evaluation Datasets
For evaluation, we assess models trained with and without DPO on three widely used benchmarks. VideoMME [8] serves as a comprehensive benchmark for general video-language understanding, MLVU [40] focuses on long-video comprehension, and Vinoground [36] specifically evaluates fine-grained temporal reasoning. To ensure high-quality video processing during inference, we increase the maximum pixels per frame to 250,000 and limit the number of frames per video to 64.
3.2 Main Results
We evaluate the effectiveness of our method by comparing the performance of three models: the original model without additional tuning, the model trained directly with supervised fine-tuning (SFT), and the model first tuned on our DPO data before undergoing SFT (Ours). The results, presented in Table 1, highlight the impact of our approach across multiple benchmarks.
Our method consistently outperforms both the original and SFT-trained models, demonstrating its ability to enhance temporal understanding in video-language models. Notably, on VideoMME, our model achieves substantial improvements on the temporal-related subtasks, with a gain of +3.6 points in Temporal Perception and +2.8 points in Temporal Reasoning. This suggests that the fine-grained alignment introduced by our DPO framework effectively reinforces the model’s capacity to process and reason about temporal information. Furthermore, our method yields a significant boost on Vinoground, a benchmark specifically designed to assess fine-grained temporal understanding, improving performance by 2.6 points—representing a relative increase of over 10%.
These results provide strong empirical evidence that our approach successfully enhances video LLMs by improving their ability to capture temporal structure. By incorporating fine-grained alignment before SFT, our method injects a more robust temporal understanding into the model, leading to consistent gains across diverse video-language tasks.
3.3 Analysis
Transferrability across models
VideoMME MLVU Vinoground
Overall Overall Text Video Group Avg
Qwen2.5-VL-7B 63.1 64.0 41.6 34.6 17.6 31.3
   + SFT 63.1 65.5 42.8 28.8 14.6 28.7
   + TEMPLE 63.8 67.3 44.2 25.8 12.4 27.5
InternVL2.5-8B 64.4 67.8 42.6 23.8 10.2 25.5
   + SFT 64.7 67.4 42.8 24.4 10.2 25.8
   + TEMPLE 64.3 67.9 45.0 24.4 10.4 26.6
Figure 5: Performance comparison between different tuning strategies. Our TEMPO brings significant boost for downstream tasks by enhancing the temporal alignment before SFT.
In our main experiment, we adopt a self-generation approach, using the same model for both generating DPO data and performing DPO. This design choice is crucial, as it ensures that the perturbations and response preferences align closely with the model’s inherent strengths and weaknesses. To assess the transferability of our DPO data across different architectures, we conduct an additional experiment where the DPO data generated by Qwen2-VL-7B is used to train two other models: Qwen2.5-VL-7B [2], which shares a similar architecture, and InternVL2.5-8B [4], which differs more significantly. The results, presented in Table 3, show that applying DPO on these models using externally generated data still yields improvements over standard SFT on certain metrics, indicating some level of transferability. However, the gains are neither as consistent nor as pronounced as those observed with self-generated data. This suggests that while DPO data can generalize to some extent, aligning the data generation process with the target model leads to more effective learning, likely due to a better match in learned representations and response tendencies.
VideoMME MLVU Vinoground
Overall Overall Text Video Group Avg
Qwen2-VL-7B 60.4 64.5 34 26.6 10.6 23.7
   + SFT 60.7 64.6 34.6 26.6 9.6 23.6
   + TEMPLE (2B-version data) 60.9 64.3 37.4 28.2 12.4 26.0
Qwen2-VL-2B 55.2 57.7 26.4 24.6 6.2 19.1
   + SFT 54.6 60.0 27.4 26.0 5.4 19.6
   + TEMPLE (7B-version data) 52.9 59.5 26.0 24.6 6.2 18.9
Figure 6: Vinoground score for models trained with DPO data of different difficulty levels.
Transferrability across model sizes
Ablation Study on Pre-SFT Alignment
To showcase the effectiveness of the Pre-SFT Alignment strategy, we conduct an ablation study comparing the performance of multiple different tuning orders. Figure 5 highlights the performance differences of Qwen2-VL-7B, comparing its results after tuning with various strategies against the original base model. The Pre-SFT alignment strategy consistently outperforms all other approaches across all three benchmarks. This demonstrates its ability to effectively combine the benefits of both realms: enhancing the model’s temporal understanding while preserving high levels of instruction-following capability. This ablation study underscores the advantages of aligning the model with fine-grained temporal knowledge before instruction fine-tuning, enabling better overall performance.
Analysis of Difficulty Scheduling
To investigate the impact of DPO data difficulty on model performance, we first conduct a preliminary comparison by tuning Qwen2-VL-7B with a single split of DPO data containing the same difficulty level and evaluating the models on Vinoground. As shown by the bars in Figure 6, models tuned with easier DPO data exhibit better performance. This suggests that, due to the initial inability of the model to capture temporal structure, preference optimization with easier data proves more beneficial in the early stages of alignment. This observation motivates our curriculum learning strategy, which formulates DPO as a staged process, progressively increasing the difficulty of DPO data throughout the course. As shown by the line in Figure 6, this progressive tuning approach leads to steady improvements in model performance, validating the efficacy of our strategy in helping the model gradually refine its temporal understanding.
4 Related Work
4.1 Video Large Languge Models
Building upon early efforts to integrate visual understanding into powerful language models [19, 18, 42], current Video LLMs typically consist of three key components: a visual encoder, a multimodal projector, and a language model [28]. The visual encoder processes video frames into patchified feature sequences, which are then transformed by the multimodal projector into discrete visual tokens. These visual tokens, along with text tokens, serve as inputs to the language model, enabling it to generate textual responses conditioned on video content. Efforts to improve video LLMs have largely followed two main directions. One approach focuses on expanding the scale and improving the quality of video datasets [39, 3, 4], often through the use of external annotation tools and proprietary LLMs to generate detailed captions. Another direction aims to extend the applicability of Video LLMs by enabling them to handle long-duration videos [37, 26, 25], high-resolution content [31, 2, 5, 13], or streaming video inputs [17]. In contrast, our work addresses a more fundamental challenge in video LLMs: the acquisition of temporal knowledge. Prior studies have frequently highlighted the deficiency of these models in capturing fine-grained temporal structure, which we directly tackle through our proposed approach.
4.2 Direct Preference Optimization
Originally introduced in the text domain as an alternative to Reinforcement Learning from Human Feedback (RLHF) [23], Direct Preference Optimization (DPO) [24] was first employed to align language model behavior with human preferences, particularly to mitigate the generation of harmful or undesirable responses. DPO optimizes an implicit reward model derived from preference data, providing a straightforward and effective approach to post-training alignment. Due to its simplicity and strong performance, DPO has gained widespread adoption, leading to various improvements such as KTO [7] and $\beta$-DPO [32]. More recently, research has explored the adaptation of DPO and other alignment strategies to multi-modal LLMs by developing carefully curated data generation pipelines. For instance, POVID [41] constructs preference data for image-language models by comparing responses to clean and noised visual inputs, while MIA-DPO [20] generates preference data for multi-image understanding by concatenating single-image datasets and assessing whether model responses correctly attend to the intended image.
Recent concurrent works [38, 15, 34] have also explored DPO for video LLMs based on model responses to corrupt visual inputs, boosting model performance on a wide range of tasks. However, our approach distinguishes itself through several key innovations. For instance, Tarsier 2 [34] applies DPO exclusively to short video clips, limiting its ability to capture long-range temporal dependencies. Additionally, both Tarsier 2 and TPO [15] are reliant on external proprietary models to curate and filter preference data. In contrast, our method introduces a carefully designed pipeline that automatically filters and selects temporally rich videos spanning 1–3 minutes and generates preference data in a fully self-sufficient manner. Furthermore, beyond dataset curation, we present two important technical innovations: difficulty scheduling, which structures training based on complexity for more effective learning, and pre-SFT alignment, which injects temporal knowledge before instruction tuning. These innovations collectively contribute to more robust and temporally-aware video reasoning, setting our approach apart from concurrent work in video DPO.
5 Conclusion
In this work, we propose a novel approach to enhance the temporal understanding of video LLMs through DPO, addressing the deficiency in capturing fine-grained temporal structures in current models. Our method leverages a carefully designed data generation pipeline to curate video preference data and a novel curriculum learning + Pre-SFT Alignment strategy to inject temporal understanding abilities into video LLMs. We demonstrate the effectiveness of our approach through extensive experiments on multiple benchmarks, showing substantial gains in temporal reasoning and perception. These results open the door for further improvements in video LLMs, particularly in refining their ability to understand and reason about temporal dynamics in videos.

### v3

@@ 3.1 Setting
We conduct our main experiments using the Qwen2-VL model family (Qwen2VL), focusing on two model scales: 2B and 7B parameters, due to its strong performance in open-source Video LLMs. We also provide additional evaluation results on Qwen2.5-VL-7B and MiMo-VL-7B in the Appendix.
Data Generation.
The videos used for preference data generation are sourced from a subset of the open-domain LLaVA-Video-178K dataset (llava-video-sft), i.e., Youtube videos within the duration of 1-3 minutes. After our data pipeline, the final data contains 6,385 preference pairs for each difficulty level, resulting in a total of 25,540 training samples.
Training Details.
For DPO training, we adopt a four-stage progressive difficulty schedule, where the model is trained sequentially on increasingly challenging data. Beginning with the easiest perturbation setting ($r=16$), the model gradually adapts to more complex data at $r=8$, $r=4$, and finally $r=2$. For SFT, we use a dataset of 60,000 samples randomly drawn from LLaVA-Video-178K (llava-video-sft). To ensure efficient training, we apply LoRA (hu2021lora) in both DPO and SFT stages. Detailed hyperparameters can be found in the Appendix.
Evaluation Datasets
For evaluation, we assess models trained with and without DPO on four widely used benchmarks. VideoMME (videomme) serves as a comprehensive benchmark for general video understanding; MLVU (MLVU) focuses on long-video comprehension; Vinoground (vinoground) and TempCompass (tempcompass) specifically evaluate fine-grained temporal reasoning. For all tasks, we use a maximum pixel per frame of 250,000 and limit the number of frames to 64.
Baselines
We validate the effectiveness of our method by comparing its performance against two baselines: the original model without additional tuning, and the model trained directly with SFT. We also report the performance of recent Video LLMs as reference points, including a range of open-source models (videollama2; timemarker; videot3; zhang2024longva; lin2023vila; li2024llava; minicpmv; lin2023video) and closed-source models (gemini; gpt4o).
3.2 Results
Table 1 summarizes the results of our main experiments. On the four benchmarks evaluated, our TEMPLE consistently improves video understanding performance compared to the original model and the model trained solely using SFT. This demonstrates that performing temporal alignment prior to SFT effectively boosts model performance, yielding greater gains compared to the SFT approach. The improvements are particularly pronounced on dimensions related to temporal understanding. On the VideoMME benchmark, TEMPLE delivers substantial gains in temporally-sensitive metrics. For instance, using the Qwen2-VL-7B model, TEMPLE achieves significant improvements over the original model: +3.6 points (70.9 → 74.5) in Temporal Perception and +2.8 points (42.4 → 45.2) in Temporal Reasoning. Furthermore, on the temporally-focused Vinoground benchmark, TEMPLE enhances performance by 3.2 points, representing a relative increase of nearly 10%. These empirical results collectively validate the effectiveness of the TEMPLE framework in enhancing models’ temporal alignment capabilities and their overall video comprehension performance, leading to consistent gains across diverse benchmarks, especially in temporal understanding tasks.
Figure 5: Performance comparison between different tuning strategies on VideoMME, MLVU, and Vinoground (Text).
Figure 6: Vinoground (Text) score for models trained with DPO data of different difficulty levels.
3.3 Analysis
Ablation Study on Pre-SFT Alignment.
To showcase the effectiveness of Pre-SFT Alignment, we conduct an ablation study comparing the performance of multiple different tuning orders. Figure 5 shows the performance differences of Qwen2-VL-7B, comparing its results after tuning with various strategies against the original base model. The Pre-SFT Alignment strategy consistently outperforms other approaches across all three benchmarks. This demonstrates its ability to effectively combine the benefits of both realms: enhancing the model’s temporal understanding while preserving high levels of instruction-following capability. Furthermore, we provide a comparison between our method against previous post-SFT video alignment methods including TPO (tpo) and LLaVA-Hound (llava-hound) in the Appendix, which demonstrates the superiority of our method and the complementarity of both strategies. This ablation study underscores the advantages of aligning the model with fine-grained temporal knowledge before instruction fine-tuning, enabling better overall performance.
Analysis of Progressive Difficulty Scheduling.
To investigate the impact of DPO data difficulty on model performance, we first conduct a preliminary comparison by tuning Qwen2-VL-7B with a single split of DPO data containing the same difficulty level and evaluating the models on Vinoground. As shown by the bars in Figure 6, models tuned with easier DPO data exhibit better performance. This suggests that, due to the initial inability of the model to capture temporal structure, preference optimization with easier data proves more beneficial in the early stages of alignment. This observation motivates our curriculum learning strategy, which formulates DPO as a staged process, progressively increasing the difficulty of DPO data throughout the course. As shown by the line in Figure 6, this progressive tuning approach leads to steady improvements in model performance, validating the efficacy of our strategy in helping the model gradually refine its temporal alignment.
Video ML Vinoground
MME VU Text Video Group
Qwen2.5-VL-7B 63.1 64.0 41.6 34.6 17.6
   + SFT 63.1 65.5 42.8 28.8 14.6
   + TEMPLE 63.8 67.3 44.2 25.8 12.4
InternVL2.5-8B 64.4 67.8 42.6 23.8 10.2
   + SFT 64.7 67.4 42.8 24.4 10.2
   + TEMPLE 64.3 67.9 45.0 24.4 10.4
Video ML Vinoground
MME VU Text Video Group
Qwen2-VL-7B 60.4 64.5 34.0 26.6 10.6
   + SFT 60.7 64.6 34.6 26.6 9.6
   + TEMPLE 2B 60.9 64.3 37.4 28.2 12.4
Qwen2-VL-2B 55.2 57.7 26.4 24.6 6.2
   + SFT 54.6 60.0 27.4 26.0 5.4
   + TEMPLE 7B 52.9 59.5 26.0 24.6 6.2
Transferrability across Model Architectures.
In our main experiment, we adopt a self-generation approach, using the same model for both generating DPO data and performing DPO. To assess the transferrability of our DPO data across different architectures, we conduct an additional experiment where the DPO data generated by Qwen2-VL-7B is used to train two other models: Qwen2.5-VL-7B (Qwen2.5VL) and InternVL2.5-8B (internvl2.5). The results, presented in Table 2, show that applying DPO on these models using externally generated data still yields improvements over standard SFT on certain metrics, indicating some level of transferability. However, the gains are neither as consistent nor as pronounced as those observed with self-generated data. This suggests that while DPO data can generalize to some extent, aligning the data generation process with the target model leads to more effective learning.
Transferrability across Model Sizes.
4 Related Work
Video Large Languge Models
builds upon early efforts to integrate visual understanding into powerful language models (liu2023llava; liu2023llava15; zhu2023minigpt4), typically consisting of three key components: a visual encoder, a multimodal projector, and a language model (video-llm-survey). This design enables the model to perform video understanding by generating textual responses conditioned on the video input. Efforts to improve Video LLMs include expanding the scale and improving the quality of video datasets (llava-video-sft; chen2024sharegpt4video; internvl2.5; DBLP:journals/corr/abs-2506-03569), and extending the applicability of Video LLMs by enabling them to handle long-duration videos (zhang2024longva; timechat; ren-etal-2023-testa), high-resolution content (Qwen2VL; Qwen2.5VL; minicpmv; li2024llava), or streaming video inputs (streamingbench). In contrast, our work addresses a more fundamental challenge in Video LLMs: the acquisition of temporal knowledge. Prior studies have frequently highlighted the deficiency of these models in capturing fine-grained temporal structure, which we directly tackle through our proposed approach.
Direct Preference Optimization
was originally introduced in the text domain to align language model behavior with human preferences. By optimizing an implicit reward model derived from preference data, DPO provides an effective approach to post-training alignment. More recently, research (povid; MIA-DPO; li-etal-2024-vlfeedback; vlrlhf) has adapted DPO and other alignment strategies to multi-modal LLMs by developing carefully curated data generation pipelines. In the video domain, some works (llava-hound; tpo; tarsier2) have also explored DPO for Video LLMs based on model responses to corrupt visual inputs. However, our approach distinguishes itself through several key improvements. First, our method introduces a carefully designed pipeline that automatically filters temporally rich videos and generates preference data in a fully self-sufficient manner, without relying on external proprietary models. Furthermore, beyond dataset curation, we present a novel Progressive Pre-SFT Alignment strategy for more effective temporal alignment, setting our approach apart from concurrent work in video DPO.
5 Conclusion
In this work, we propose a novel framework to enhance the temporal alignment of Video LLMs through DPO, addressing the lack of supervision signals in current training paradigms. Our method leverages a carefully designed data generation pipeline to curate temporality-oriented video preference data and a novel Progressive Pre-SFT Alignment strategy to inject temporal understanding abilities into Video LLMs. We demonstrate the effectiveness of our approach through extensive experiments on multiple benchmarks, showing substantial gains in temporal understanding. These results open the door for further improvements in Video LLMs, particularly in refining their ability to understand and reason about temporal dynamics in videos.
@@ 5 Conclusion
Table 4: The number of remaining videos after each step of our video selection pipeline.
Model VM ML LVB ES PT Vinoground TempCompass
V T G CM C MC YN
MiMo-VL-7B+SFT 63.52 65.38 58.27 57.80 65.71 41.2 29.2 14.8 82.43 54.14 73.54 73.09
MiMo-VL-7B+TEMPLE 63.26 65.30 58.56 58.60 65.81 43.6 30.0 16.8 82.30 60.08 73.29 73.18
Qwen2.5-VL-7B+SFT 63.15 66.43 57.89 55.40 69.50 43.8 27.2 12.6 81.84 53.69 73.67 74.56
Qwen2.5-VL-7B+TEMPLE 63.48 67.11 58.56 53.80 69.01 44.0 29.2 15.0 82.10 55.34 73.29 74.56
Table 5: Performance of TEMPLE using Qwen2.5-VL-7B-Instruct and MiMo-VL-7B-SFT as base models.
Model VideoMME MLVU LongVideoBench EgoSchema PerceptionTest Average
SFT 60.67 64.60 55.12 50.00 60.09 58.10
TEMPLE-SFT 60.96 65.79 55.35 52.60 60.82 59.10 (+1.00)
SFT-TPO 60.19 63.93 56.62 52.60 60.78 58.82 (+0.72)
SFT-Hound 60.30 63.92 55.80 51.40 60.56 58.40 (+0.30)
TEMPLE-SFT-TPO 60.70 64.63 56.40 54.80 61.49 59.60 (+1.50)
TEMPLE-SFT-Hound 60.89 64.54 56.02 53.40 61.10 59.19 (+1.09)
Table 6: Performance of different combinations of TEMPLE and post-SFT DPO methods on general video understanding benchmarks.
Model Vinoground TempCompass Average
Video Text Group Matching Captioning Multi-Choice Yes/No
SFT 34.6 26.6 9.6 80.51 46.71 70.82 74.84 45.91
TEMPLE-SFT 37.2 29.6 12.2 79.51 47.65 71.39 74.60 47.31 (+1.40)
SFT-TPO 34.6 25.4 9.8 80.31 49.80 70.76 74.89 46.11 (+0.20)
SFT-Hound 33.6 28.4 9.8 80.31 49.20 70.82 74.68 46.34 (+0.42)
TEMPLE-SFT-TPO 35.6 27.2 11.2 79.51 51.50 71.33 74.40 46.93 (+1.02)
TEMPLE-SFT-Hound 36.4 29.0 11.0 79.64 50.35 71.01 74.89 47.22 (+1.31)
Table 7: Performance of different combinations of TEMPLE and post-SFT DPO methods on temporal understanding benchmarks.
Appendix A Data
For our experiments, we generate preference data using videos from LLaVA-Video-180K, a comprehensive conversation-style dataset designed for fine-tuning Video LLMs. Its broad coverage of video content and duration makes it an ideal source for our pipeline. However, we observe that a significant portion of videos in this dataset exhibit low dynamic content and insufficient temporal information, a common limitation in video datasets. Examples include static interview footage with alternating speaker shots or multi-view recordings of a single object/action.
This observation motivates our video filtering pipeline which selects temporally informative videos based on scene number and inter-scene similarity. Table 4 presents the number of remaining videos after each step of our video selection pipeline. The large number of discarded videos, especially for shorter video within the duration of 1-2 minutes, validates our observation and underscores our pipeline’s necessity in generating temporality-oriented preference data.
Appendix B Training Details
@@
Our DPO process is split into four stages where each stage corresponds to a single difficulty factor. During DPO, we apply LoRA (hu2021lora) with rank 32, $\alpha=64$, a batch size of 16, and a learning rate of 5e-6, while keeping the ViT module frozen. We adopt Adam (adam) as the optimizer throughout training. Each DPO stage lasts for approximately one hour on 8 NVIDIA H800 GPUs, accumulating to a total of four hours for the entire alignment process.
@@
During SFT, the ViT module is unfrozen to allow for additional adaptation. The SFT phase follows a similar training setup with LoRA rank set to 32, $\alpha$ set to 32, a batch size of 16, and a higher learning rate of 1e-4. The entire fine-tuning process takes approximately three hours on the same hardware setup.
@@
Both DPO and SFT follow the original Qwen2-VL preprocessing pipeline, where videos are sampled at 2 FPS. Note that while extracted keyframes are used for data construction, they are not used during training. To accommodate memory constraints, we limit the number of pixels per frame to 90,000, the maximum number of frames per video to 100, and the total sequence length to 8,192.
Appendix C Additional Results
Evaluation on additional base models.
To demonstrate the broad applicability of our method, we apply TEMPLE to Qwen2.5-VL-7B-Instruct and MiMo-VL-7B-SFT and evaluate model performance on an extended set of benchmarks including VideoMME (VM), MLVU (ML), LongVideoBench (LVB), EgoSchema (ES), PerceptionTest (PT), Vinoground (Text, Video, Group), and TempCompass (Caption Matching, Captioning, Multi-Choice, Yes/No) (videomme; MLVU; LongVideoBench; mangalam2023egoschema; perceptiontest; vinoground; tempcompass). The results are presented in Table 5. Results show that TEMPLE consistently enhances performance over SFT across a majority of the evaluated benchmarks. Notably, for the Vinoground benchmark, TEMPLE improves all three metrics for both base models, highlighting its strength in temporal alignment. Beyond this, TEMPLE also demonstrates gains in general video understanding benchmarks. These improvements, observed across two distinct base models, underscore the generalizability and effectiveness of TEMPLE for enhancing video-language models.
Comparison with other video DPO methods.
We further compare our Pre-SFT Alignment strategy with other video DPO methods which follow the conventional post-SFT DPO strategy. Specifically, we perform TPO (tpo) and LLaVA-Hound-DPO (llava-hound) after training the model with SFT. During this post-training process, we adopt the same hyper-parameters as our DPO process in TEMPLE. We also experimented with combining TEMPLE with post-SFT DPO methods to see how these two strategies work together. The results on general video understanding and temporal understanding benchmarks are presented in Table 6 and Table 7, respectively.
These results lead to several key observations. First, all evaluated methods, including TEMPLE, TPO, and LLaVA-Hound-DPO, consistently outperform the SFT baseline across both benchmark categories. Second, TEMPLE alone surpasses both TPO and LLaVA-Hound-DPO when applied individually, achieving higher average scores on general video understanding (59.10 vs. 58.82 for TPO and 58.40 for Hound) and temporal understanding benchmarks (47.31 vs. 46.11 and 46.34, respectively), underscoring the advantage of pre-SFT temporal alignment. Third, the two approaches exhibit complementary strengths: TPO and Hound excel in specific tasks such as LongVideoBench and TempCompass, while TEMPLE demonstrates superior performance on others like Vinoground. The orthogonality of these strategies is affirmed by the additional gains observed when combining them, which promotes the joint application of both paradigms for comprehensive performance improvement.
Acknowledgments
This research was partially supported by the National Natural Science Foundation of China under Grant No. 92470205 and No. 62176002. Xu Sun is the corresponding author.